# Synthetic Geology - Structural Geology Meets Deep Learning

### v1

@@ Abstract
Visualizing the first few kilometers of the Earth’s subsurface, a long-standing challenge gating a virtually inexhaustible list of important applications, is coming within reach through deep learning. Building on techniques of generative artificial intelligence applied to voxelated images, we demonstrate a method that extends surface geological data supplemented by boreholes to a three-dimensional subsurface region by training a neural network. The Earth’s land area having been extensively mapped for geological features, the bottleneck of this or any related technique is the availability of data below the surface. Drilling is expensive and generally confined to regions where recovery of minerals, petroleum, water, thermal energy, or other marketable commodities is economical. We close this data gap in the development of subsurface deep learning by designing a synthetic data-generator process that mimics eons of geological activity such as sediment compaction, volcanic intrusion, and tectonic dynamics to produce a virtually limitless number of samples of the near lithosphere. A foundation model trained on such synthetic data is able to generate a 3D image of the subsurface from a previously unseen map of surface topography and geology, showing increasing fidelity with increasing access to borehole data, depicting such structures as layers, faults, folds, dikes, and sills. We illustrate the early promise of the combination of a synthetic lithospheric generator with a trained neural network model using generative flow matching. Ultimately, such models will be fine-tuned on data from applicable campaigns, such as mineral prospecting in a given region. Though useful in itself, a regionally fine-tuned models may be employed not as an end but as a means: as an AI-based regularizer in a more traditional inverse problem application, in which the objective function represents the mismatch of additional data with physical models with applications in resource exploration, hazard assessment, and geotechnical engineering.
Introduction
@@
Structural geology focuses on the deformation and motion of rocks in the Earth’s crust, describing three-dimensional distributions of rock formations and structures to understand the processes that shape the crust [1, 2]. Structural geologists use a variety of techniques and methods to investigate and interpret these features, including field mapping, structural analysis of rock outcrops, laboratory experiments, remote sensing, and geophysical imaging. The insights gained are essential for understanding the evolution of mountain ranges, the formation of mineral and energy resources, and the assessment of geological hazards such as earthquakes and landslides, among other fundamental aspects of Earth science.
@@
The construction of accurate structural geological models from collected data is a complex, ill-posed problem fraught with challenges (see, e.g., [3] and references within). These models must reconcile geological observations obtained from what are often sparse and incomplete sources such as field data, rock samples, geophysical surveys, and remote sensing. Geological data is subject to uncertainty and interpretation, posing significant challenges in quantifying and addressing errors within the model [4]. The interpretation of uncertain data is further exacerbated by the intricate and heterogeneous nature of geological structures, which unfold over diverse spatial and temporal scales– ranging from microscopic mineral deformation to regional tectonic movements spanning millions of years [5]. Integrating disparate datasets and models from multiple disciplines, including structural geology, geophysics, and geochemistry, represents an additional hurdle that demands interdisciplinary collaboration [6].
@@
Overcoming these challenges requires innovative technologies, advanced analytical methods, and interdisciplinary collaboration. However, the rewards of overcoming these obstacles are substantial, as accurate and reliable structural geological models are indispensable for understanding Earth’s subsurface processes, predicting geological hazards, and managing natural resources. Traditionally, the construction of 3D structural geological models has been a manual, labor-intensive process, relying on expert interpretation, visualization, and iterative refinement of multiple data sources [7, 8, 9]. Such manual workflows can take many hours and usually produce a single deterministic model, ignoring the underlying ill-posedness of the problem.
In this study, we take a step toward automating structural geology modeling by leveraging generative artificial intelligence (GAI) techniques. First, we create a geologically plausible, infinite synthetic structural dataset that reflects the diverse processes and events shaping the Earth’s crust. This dataset is then used to train a 3D GAI model with flow matching [10, 11]. Unlike traditional deterministic methods, this approach allows for the generation of multiple plausible models, capturing a richer range of potential subsurface scenarios.
@@
Once trained, the generative model can be integrated with other conditional data and applied to various geological tasks. For example, it can be coupled with information obtained from drilling or surface measurement, geostatistical interpolation, or geological segmentation that incorporates geological priors, improving model reliability and reducing ambiguity in subsurface interpretations. We demonstrate this capability through experiments on the reconstruction of geological stratigraphy from sparse geological data, illustrating the potential of generative models to advance structural geology research and practice.
Methods
Synthetic Dataset
@@ Methods
Recent advances in generative diffusion models capable of strikingly realistic images (see, e.g., [12, 11] and references within) have both motivated and been fueled by the compilation of ever-expanding, detailed training datasets (e.g., Imagenet [13]). However, for 3D geological data, no comparable large-scale training sets exist, partly because we cannot directly observe high-resolution, three-dimensional underground rock formations. Consequently, one of the primary obstacles to creating a geological training set for 3D GAI is acquiring comprehensive samples that reflect the wide variety of possible features and configurations that are present in the Earth’s crust [14]. To address this gap in information, we generate a synthetic dataset that captures the fundamental principles of structural geology that have been amassed through research and field observation. The ideal synthetic data for deep learning should simulate empirically drawn, independent, and identically distributed (i.i.d.) samples from the set of all 3D volumes in the Earth’s crust. A deep generative model aims to learn a representation of the underlying probability distribution of geological structures from the samples, either implicitly or explicitly, in order to generate new samples that have never been seen before [15, 16].
@@
The difficulty of obtaining true observational data extends beyond geology, affecting other domains in scientific machine learning; fluid flow [17], wave simulation [18], and biology [19] share similar challenges. Moreover, progressively more complex large language models are predicted to run out of sufficient training data in the future [20]. In these cases, it has been suggested that numerical simulations or synthetic data are used to either generate or augment the data that is needed to train a generative model. These simulations can produce abundant, high-resolution data at large scales, yet also carry the risk that the generated data has some implicit bias and does not correctly represent the true distribution. Nevertheless, such theoretical frameworks represent our best current knowledge about the particular field of interest, making them highly useful for explaining indirect measurements.
@@
In response to these constraints, we adopt a geomodel-driven approach [4], bootstrapping an initial synthetic dataset through classical simulation methods based upon fundamental principles of how stratigraphic layers form over long time spans. We compile sequential, parameterized geological processes and workflows via mesh- and grid-based numerical techniques. These simulations are frequently used to explain observed geological formations, making them well-suited for creating synthetic data.
Current open-source 3D modeling efforts have been led by tools such as GemPy [21], Noddy [22, 23], and LoopStructural [24], producing 3D geological volumes using a set of informed and constrained parameters to model limited underlying geological processes over time. While these tools are invaluable in their respective scopes of use, they do not natively support large-scale randomization. For example, Jessell et al. [23] have suggested that more parameters, more events, linked events, and more models would improve their existing efforts. We thus adopt a computational method suggested in Visible Geology [25], where geological processes are represented as mathematical transformations of a 3D mesh. By combining multiple transformations, one can generate a wide range of realistic geological models. In fact, such models are used to train geologists [25], suggesting their suitability as a synthetic dataset for training machine-learning models. Additionally, randomizing and reordering these processes enables the generation of a virtually limitless number of 3D volumes depicting complex and varied geology. We implement this randomized workflow in the open-source Python package StructuralGeo.
Sample Generation
@@
Geological model training data samples, referred to as Geomodels, take the form of 3D tensors denoted as ${\bf m}({\bf x}):R^{3}\rightarrow{\mathbb{N}},\ \ {\bf x}=[{\bf x}_{1},{\bf x}_{2},{\bf x}_{3}]$ and =${\mathbb{N}}=\{1,\ldots,N\}$. Each element of the tensor ${\bf x}$ corresponds to a voxel in a 3D mesh grid, with an integer label (from $1,\ldots,N$) indicating its assigned rock-strata category. The data is composed of discrete categorical values, which are representative of the labeling and segmentation that a geologist would apply to collected field data.
@@
    Transformations $\mathcal{T}:R^{3}\rightarrow R^{3}$ which transport or displace the geological medium, and
@@
    Depositions $\mathcal{D}:R^{3}\rightarrow\mathbb{N}$ which modify the composition.
Transformations:
Transformations $(\mathcal{T})$ are general operations that deform, rotate, or otherwise displace the geological medium according to parameters ${\bm{\theta}}$. Let ${\bf u}({\bf x};{\bm{\theta}})\colon R^{3}\to R^{3}$ be a vector-valued function dependent on ${\bm{\theta}}$. A single transformation ${\cal T}_{{\bm{\theta}}}$ is then expressed as a coordinate mapping, that is
$\mathcal{T}_{{\bm{\theta}}}\bigl{(}{\bf m}({\bf x})\bigr{)}={\bf m}\bigl{(}{\bf x}+{\bf u}({\bf x};{\bm{\theta}})\bigr{)}.$ (1)
@@
Depositions:
Depositions $(\mathcal{D})$ add or remove material into the model domain through sedimentation, erosion, or intrusion events; for example, the transport of hot fluids from the interior of the Earth through a fissure. In a deposition, we envision some material (new rock) that replaces existing rock by pushing it aside, eroding it away, or accumulating above it. The extent and nature of a deposition event are controlled by a set of parameters ${\bm{\theta}}$. A simple mathematical model uses a window function ${\bf w}({\bf x},{\bm{\theta}})$ that has the value $1$ where the deposition occurs and 00 otherwise. This function is generated by first considering a canonical window
${\bf w}_{0}({\bf x})=\prod_{i=1}^{3}(H({\bf x}_{i}+1)-H({\bf x}_{i}-1)),$
@@
To numerically evaluate the model ${\bf m}({\bf x})$ under transformations and depositions, we use backward interpolation, common in image transformation problems [26]. Figure 1 illustrates one example of such a history.
Although the Geomodel equation 4 is deterministic once the sequence and parameters are fixed, we introduce variability using a Markov chain to sequence new histories, drawing from the set of all possible geological processes $\{{\cal P}^{j}\},j=1,\ldots,k$. After each process, the next is chosen with some probability
@@
conditioned on the previous event. This yields a random geological history for the computational sequence in equation 4. Termination occurs when a special end-of-process token is drawn; hence, models emerge as random computational sequences of variable length. Because each process has its randomized parameters, repeated runs produce diverse realizations with high variability in both the ordering of events and overall geological age. Figure 2 provides an overview of the generation pipeline.

### v3

# Synthetic Geology: Structural Geology Meets Deep Learning

@@ Abstract
Reconstructing the structural geology and mineral composition of the first few kilometers of the Earth’s subsurface from sparse or indirect surface observations remains a long-standing challenge with critical applications in mineral exploration, geohazard assessment, and geotechnical engineering. This inherently ill-posed problem is often addressed by classical geophysical inversion methods, which typically yield a single maximum-likelihood model that fails to capture the full range of plausible geology. The adoption of modern deep learning methods has been limited by the lack of large 3D training datasets. We address this gap with StructuralGeo, a geological simulation engine that mimics eons of tectonic, magmatic, and sedimentary processes to generate a virtually limitless supply of realistic synthetic 3D lithological models. Using this dataset, we train both unconditional and conditional generative flow-matching models with a 3D attention U-net architecture. The resulting foundation model can reconstruct multiple plausible 3D scenarios from surface topography and sparse borehole data, depicting structures such as layers, faults, folds, and dikes. By sampling many reconstructions from the same observations, we introduce a probabilistic framework for estimating the size and extent of subsurface features. While the realism of the output is bounded by the fidelity of the training data to true geology, this combination of simulation and generative AI functions offers a flexible prior for probabilistic modeling, regional fine-tuning, and use as an AI-based regularizer in traditional geophysical inversion workflows.
[…]
JGR: Machine Learning and Computation
Department of Computer Science, University of British Columbia Department of Earth, Ocean and Atmospheric Sciences, University of British Columbia Division of Computer, Electrical and Mathematical Sciences and Engineering, King Abdullah University of Science and Technology
[…]
Simon Ghyselinckssghyseli@cs.ubc.ca
{keypoints}
Synthetic geology enables training of deep learning models for geophysics.
Generative flow matching reconstructs 3D geology from surface maps and sparse borehole data, giving multiple probable scenarios.
Supports geophysical inversions, probabilistic modeling, and resource exploration.
Plain Language Summary
Understanding and reconstructing the layers and structures of rock beneath the Earth’s surface is an important but challenging task for geologists. It has practical uses in assessing geohazards such as earthquakes and landslides, and in guiding decisions for mineral and oil exploration. Most of our information for reconstruction comes from surface measurements, since direct observations are available only from narrow drill cores or mines. New and exciting artificial intelligence (AI) tools that can create realistic images or animations could be adapted to geology, but they require large swaths of training data which we do not have. In this study, we develop StructuralGeo, a geological computer simulation that reproduces geological processes over millions of years to create realistic examples of what lies hidden underground. The simulated data is used to train a generative AI that produces multiple 3D scenarios from existing surface maps and drill cores. The results can be used to enhance traditional techniques, for example by estimating the possible size range of an underground rock body based on limited drilling samples.
1 Introduction
@@
Structural geology focuses on the deformation and motion of rocks in the Earth’s crust, describing three-dimensional distributions of rock formations and structures to understand the processes that shape the crust [fossen2016structural]. Structural geologists use a variety of techniques and methods to investigate and interpret these features, including field mapping, structural analysis of rock outcrops, laboratory experiments, remote sensing, and geophysical imaging. The insights gained are essential for understanding the evolution of mountain ranges, the formation of mineral and energy resources, and the assessment of geological hazards such as earthquakes and landslides, among other fundamental aspects of Earth science.
@@
The construction of accurate structural geological models from collected data is a complex, ill-posed problem fraught with challenges [mcclay2013mapping]. These models must reconcile geological observations obtained from what are often sparse and incomplete sources such as field data, rock samples, geophysical surveys, and remote sensing. Geological data is subject to uncertainty and interpretation, posing significant challenges in quantifying and addressing errors within the model [wellmann2018ChapterOne3Da]. The interpretation of uncertain data is further exacerbated by the intricate and heterogeneous nature of geological structures, which unfold over diverse spatial and temporal scales ranging from microscopic mineral deformation to regional tectonic movements spanning millions of years [karpatne2019MachineLearningGeosciencesa]. Integrating disparate datasets and models from multiple disciplines, including structural geology, geophysics, and geochemistry, represents an additional hurdle that demands interdisciplinary collaboration [wellmann2010towards].
@@
Overcoming these challenges requires innovative technologies, advanced analytical methods, and interdisciplinary collaboration. However, the rewards of overcoming these obstacles are substantial, as accurate and reliable structural geological models are indispensable for understanding Earth’s subsurface processes, predicting geological hazards, and managing natural resources. Traditionally, the construction of 3D structural geological models has been a manual, labor-intensive process, relying on expert interpretation, visualization, and iterative refinement of multiple data sources [calcagno2008geological, houlding20123d]. Such manual workflows can take many hours and usually produce a single deterministic model, ignoring the underlying ill-posedness of the problem, [R2-C2-2]
In this study, we take a step toward automating inverse modeling[R2-C2-3] of structural geology from sparse or indirect observations (e.g., boreholes)[R2-C2-4] using generative artificial intelligence (GAI). Our goal is to develop a framework capable of parsing collected field data into coherent 3D geological reconstructions. Specifically, we aim to generate multiple plausible 3D models of lithology that reflect the range of geological scenarios consistent with a given set of observations. To achieve this, we combine geological simulation with a deep generative learning framework.[R2-C3-2]
First, in Sections 2.1 and 2.2, we develop StructuralGeo [StructuralGeo_v1], a stochastic geology simulator that rapidly produces diverse 3D lithological volumes by composing representative tectonic, magmatic, and sedimentary events. Next, in Section 2.3, we describe training a 3D flow-matching model [albergoStochasticInterpolantsUnifying2023, lipman2023flow] fed by on-demand samples from StructuralGeo. Flow matching regresses a continuous-time velocity field and is closely related to diffusion models [yang2022diffusion], providing a method of indirectly sampling from the distribution of simulated synthetic geology.
@@
Once trained, the generative model, unlike the geology simulator, can be integrated with conditional data and applied to various geological tasks, described further in Section 2.4. For example, it can be coupled with information obtained from drilling or surface measurement, geostatistical interpolation, or geological segmentation that incorporates geological priors, improving model reliability and reducing ambiguity in subsurface interpretations. We demonstrate this capability in Section 3 through experiments on the reconstruction of geological stratigraphy from borehole and surface lithology, illustrating the potential of generative models to advance structural geology research and practice.
2 Methods
2.1 Synthetic Dataset
@@
Recent advances in generative diffusion models capable of strikingly realistic images [yang2022diffusion] has been fueled by the availability of increasingly large, detailed training datasets such as Imagenet [imagenet_cvpr09]. However, for 3D geological data, no comparable large-scale training sets exist, partly because we cannot directly observe high-resolution, three-dimensional underground rock formations. The primary obstacle is acquiring a comprehensive set of samples that reflects the wide variety of features and configurations that are present in the Earth’s crust [hadid2024WhenGeoscienceMeets]. To address this gap in information, we generate a synthetic dataset that captures the fundamental principles of structural geology that have been amassed through research and field observation. The ideal synthetic data for deep learning should simulate empirically drawn, independent, and identically distributed (i.i.d.) samples from the set of all 3D volumes in the Earth’s crust. A deep generative model aims to learn a representation of the underlying probability distribution of geological structures from the samples, either implicitly or explicitly, in order to generate new samples that have never been seen before [ruthotto2021IntroductionDeepGenerative].
@@
The difficulty of obtaining true observational data extends beyond geology, affecting other domains in scientific machine learning; fluid flow [lienen2024zero] and biology [goshisht2024machine] share similar challenges, for example. Moreover, progressively more complex large language models are predicted to run out of sufficient training data in the future [Jones2024]. In these cases, it has been suggested that numerical simulations or synthetic data are used to either generate or augment the data that is needed to train a generative model. Although such simulations can generate abundant, high-resolution data, they inevitably carry implicit biases that reflect the assumptions and simplifications of the underlying geological modeling. The realism of the simulated samples is ultimately bounded by those assumptions, which may introduce structural biases into any downstream inference. This limitation is not unique to synthetic data generation, as existing regularization and Bayesian inference methods likewise encode priors that shape and potentially bias the solution space [somersallo, Ch. 3.3]. Despite this challenge, such theoretical frameworks represent our best current knowledge about the particular field of interest, making them highly useful for explaining[R2-C3-4] indirect measurements [R2-C3-4]. We view the framework presented ahead as a step toward more specialized, data-driven priors[R2-C3-5] that can evolve with higher quality simulation[R2-C3-5] and empirically evaluated simulation parameters.
@@
In response to these constraints, we adopt a Geomodel-driven approach [wellmann2018ChapterOne3Da], bootstrapping an initial synthetic dataset through classical simulation methods based upon fundamental principles of how stratigraphic layers form over long time spans. We compile sequential, parameterized geological processes and workflows via mesh- and grid-based numerical techniques. This type of simulation is frequently used to explain observed geological formations.
[…]
2.2 Sample Generation
[…]
@@ 2.2 Sample Generation
wellmann2018ChapterOne3Da,[R2-C2-1] Geological model training data samples, referred to as Geomodels take the form of 3D tensors denoted as ${\bf m}({\bf x}):\mathbb{R}^{3}\rightarrow{\mathbb{N}},\ \ {\bf x}=[{\bf x}_{1},{\bf x}_{2},{\bf x}_{3}]$ and ${\mathbb{N}}=\{1,\ldots,N\}$. Each element of the tensor $\bf x$} corresponds to a voxel in a 3D mesh grid, with an integer label (from $1,\ldots,N$) indicating its assigned rock-strata category. The data is composed of discrete categorical values, which are representative of the labeling and segmentation that a geologist would apply to collected field data[R2-C2-5].
@@
    Transformations $\mathcal{T}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ which transport or displace the geological medium, and
@@
    Depositions $\mathcal{D}:\mathbb{R}^{3}\rightarrow\mathbb{N}$ which modify the composition.
2.2.1 Transformations
Transformations $(\mathcal{T})$ are general operations that deform, rotate, or otherwise displace the geological medium according to parameters $\bm{\theta}$}. Let ${\bf u}({\bf x};{\bm{\theta}})\colon\mathbb{R}^{3}\to\mathbb{R}^{3}$ be a vector-valued function dependent on $\bm{\theta}$}. A single transformation ${\cal T}_{{\bm{\theta}}}$ is then expressed as a coordinate mapping, that is
$\mathcal{T}_{{\bm{\theta}}}\bigl({\bf m}({\bf x})\bigr):={\bf m}\bigl({\bf x}+{\bf u}({\bf x};{\bm{\theta}})\bigr).$ (1)
@@ 2.2.1 Transformations
2.2.2 Depositions
Depositions $(\mathcal{D})$ add or remove material into the model domain through sedimentation, erosion, or intrusion events; for example, the transport of hot fluids from the interior of the Earth through a fissure. In a deposition, we envision some material (new rock) that replaces existing rock by pushing it aside, eroding it away, or accumulating above it. The extent and nature of a deposition event are controlled by a set of parameters $\bm{\theta}$}. A simple mathematical model uses a window function ${\bf w}({\bf x},{\bm{\theta}})$ that has the value  $1$ where the deposition occurs and 0 otherwise. This function is generated by first considering a canonical window
${\bf w}_{0}({\bf x}):=\prod_{i=1}^{3}(H({\bf x}_{i}+1)-H({\bf x}_{i}-1)),$
@@ 2.2.3 Geological Histories
To numerically evaluate the model ${\bf m}({\bf x})$ under transformations and depositions, we use backward interpolation, common in image transformation problems [Modersitzki2004]. Figure 1 illustrates one example of such a history.
Although the Geomodel (4) is deterministic once the sequence and parameters are fixed, we introduce variability using a Markov chain to sequence new histories, drawing from the set of all possible geological processes $\{{\cal P}^{j}\},j=1,\ldots,k$. After each process, the next is chosen with some probability
@@
conditioned on the previous event. This yields a random geological history for the computational sequence in (4). Termination occurs when a special end-of-process token is drawn; hence, models emerge as random computational sequences of variable length. Because each process has its randomized parameters, repeated runs produce diverse realizations with high variability in both the ordering of events and overall geological age. Figure 2 provides an overview of the generation pipeline.
Figure 1: Example of geological history illustrating the application of transformations and depositions. Panels: (A) initial strata ${\bf m}_{0}({\bf x})$, (B) fold $\mathcal{T}_{\text{fold}}$, (C) dike $\mathcal{D}_{\text{dike}}$, and (D) fault $\mathcal{T}_{\text{fault}}$. Processes $\mathcal{P}_{i}$ are sequenced randomly from a set ${\cal P}_{j},j=1,\ldots,k$ using a Markov chain, leading to varied outcomes.
Figure 2: Pipeline for random geological model generation. A batch request triggers a Markov chain sampler which selects and sequences geological processes (e.g., folding, deposition) from a transition matrix; parameters such as amplitude and wavelength are drawn from tuned random variables; the sequences are applied to produce batches of 3D tensors.
2.2.4 Limitations
The above simulation can generate a virtually infinite number of geological models with variable complexity and at any specified resolution, which can be used as data for a generative model. Nonetheless, just like any other simulation, it depends on a number of hyperparameters. Every geological process depends on a set of parameters that are sampled as random variables and must be tuned such that the outcome is “realistic” as visually determined by geologists. In addition, the transition probabilities $\pi_{j}$ between processes must be chosen. As part of development , representative samples were shown to domain experts to verify the large-scale geometries produced by the simulator, though feedback has been strictly qualitative and informal. However, it is important to realize that such models represent what experts consider to be unbiased geology. Since there are no direct observations on the geology at the (3D) resolution and scale that we need, it is virtually impossible to know if such models introduce implicit bias. Nonetheless, such models are extremely useful in generating what experts believe is a realistic Earth model.
While multiscale and multifractal variability are well established characteristics of natural geological and geophysical fields (e.g., [Dolan98, li2003scale, gagnon2003, lovejoy2007]), the present simulations target a large scale regime and[R2-C4-1] with discrete lithological categories. The simulation dataset is configured to capture characteristics in the $30-60\,\text{m}^{3}$ resolution range, emphasizing the large-scale geometry and spatial arrangement of structural domains rather than fine-scale heterogeneity. The segmentation of geology into discrete rock types serves to remove variation in favor of large-scale classification.[R2-C4-1] The current implementation of StructuralGeo is constrained to applications where this type of variability is not desired or required.
2.3 Synthetic Geology Using Generative AI
We define the simulator’s output distribution of Geomodels over all possible 3D volumes as $\pi_{\bf m}$, and aim to train a GAI model that can approximate and draw new samples from this distribution[R2-C3-3] without relying on the original simulation engine. The simulator provides the training data, and the GAI learns to produce new examples from $\pi_{\bf m}$ in a probabilistic manner, analogous to training an image generator on a large dataset of photos. However, simulator outputs have the advantage of being computed and tuned on demand, providing endless variation with a low memory storage footprint compared to fixed datasets.
Let ${\bf m}({\bf x})$ be a geological model generated by the simulation process using a sampled geological history ${\cal H}({\bf m}_{0})$, see (4). Given a method to generate virtually unlimited samples ${\bf m}({\bf x})\!\sim\!\pi_{\bf m}$, we now have the tools to train a foundation model that indirectly learns the underlying probability characteristics of any given Geomodel, $\pi_{{\bf m}}({\bf m})$. For example by learning its associated score function $s_{\bf m}:=-{\bm{\nabla}\,}_{\bf m}\log(\pi_{\bf m}({\bf m}))$ [song2019generative], or, as in our case, through the related flow-matching objective described in Section 2.3.1. Note that $\bf m$} represents discrete categorical data, so $\pi_{\bf m}$ is not directly differentiable with respect to $\bf m$}. We address this by adopting a continuous embedding and decoding scheme that enables differentiation in the model’s latent space.
While the original simulation process can generate diverse Geomodels, it is hard, and in some cases virtually impossible, to use the simulation for any other application besides the forward process of generating a random 3D geological volume from a parameterized history ${\cal H}({\bf m}_{0})$. In contrast, foundation models support a broader range of applications that vary from generation to conditional generation or inversion, enabling the incorporation of other data or constraints based on conditional distributions [chung2022come].
For instance, while the simulation process requires a structured history of geological events, a foundation model can generate realistic samples conditioned directly on partial observations in the spatial domain $\bf x$}, such as borehole data or surface measurements. Since our ultimate goal is to fuse multiple sources of geoscientific data, we seek to replace the rigid simulation process with a flexible, trainable foundation model.